A male patient with a BMI 39 and height of 66 in suffered a stroke and required mechanical ventilation. Develop a tube feeding regimen.

This problem does not have body weight.
Answer #1

Height: 66 in = 1.68 m
BMI: 39
BMI = Weight (kg) ÷ height ^2 (m^2)
Weight: BMI × Height ^2 = 39× 1.68^2 = 39 × 2.82 = 110 kg
Tube feeding regimen:
The daily amount of calories required by the patient is calculated using Harris-Benedict formula. The first step is to calculate the basal metabolic rate.
BMR= 10 × Weight + 6.25 × Height- 5 × age + 5
Then the next step is to calculate the calorie requirement. For this the BMR is multiplied with the activity factor. For sedentary: BMR × 1.2, Light activity: BMR× 1.375, moderate activity: BMR × 1.55, Very active: BMR × 1.725
For water requirements: 35 ml / kg